# The EMPI Code for Plasma-Induced Effects on Radio Waves I: Non-Magnetized Media and Applications to Fast Radio Bursts

### v1

@@ 4.1.5 Resolution Test
The effect of resolution on the received signal is demonstrated in Figure 14. With a fixed plasma screen and observer position, the resolution is varied by adjusting the grid size from $1$ AU down to $10^{-5}$ AU. At coarser resolutions, the results exhibit significant instability and large fluctuations in the signal. At such resolutions, some frequencies may exhibit signal arrivals while others do not, resulting in considerable variability. Additionally, regions where light reaches the observer tend to display similar intensities because only a limited number of rays contribute to the signal.
As the resolution is refined, the results gradually converge toward a stable state. With finer grids, the simulated waterfall plots become sharper and the signal intensity more concentrated. Notably, the frequency range over which the caustics align near the observer increases in intensity. This convergence demonstrates that improving the resolution leads to a more accurate representation of the light distribution and a more faithful reproduction of the true plasma lensing effect.
@@ 4.2 Extended Results
Building on the foundational cases discussed previously, this section present scenarios that incorporate conditions more representative of real-world observations. These examples offer a more nuanced view of the applicability of the code to complex astrophysical phenomena, demonstrating its adaptability to intricate spatial configurations and variable parameters.
@@
Figure 15 displays the intensity distribution on the observer plane for a two-dimensional Gaussian lens. The lens is modeled by the distribution
@@
which is analogous to the one-dimensional Gaussian case (with $N_{\rm e}=n_{\rm e}l$, where $n_{\rm e}$, $l$, and $a$ defined as in the 1D formulation). n this example, it sets $N_{\rm 0}=0.015\,\rm pc\,cm^{-3}$, and $a=2$ AU, with both the source-to-screen and screen-to-observer distances equal to $1$ kpc. The source is modeled as a point source with a half-open angle of $10^{-5}$ degrees, ensuring that the projected circle on the screen encompasses the entire lens. The color bar indicates the relative intensity normalized by the intensity in the absence of lensing.
Generally, the lens exerts a more pronounced impact at lower frequencies. Numerical calculations reveal bright annular regions (caustics) with significantly enhanced relative intensities, where theoretical treatments predict divergence. The lens behaves as a diverging element, redistributing light from the central region toward the periphery and thereby forming two distinct caustic rings. In regions distant from the lens, where the electron density is nearly uniform, no lensing occurs. This behavior is consistent with the findings of Clegg et al. (1998), which indicate that inner caustics are less sensitive to variations than outer caustics. The red star in Figure 15 marks the location of the observer, corresponding to the waterfall plots presented in Figure 16.
@@
Figure 16 illustrates the waterfall plots at the observer. These plots highlight how the lensing structure alters the received signal, manifesting as variations in intensity and spectrum. The DM value of the plasma lens is sufficiently low and any background DM has been removed, resulting in a flat-top square wave pattern in the waterfall plot. Although the intrinsic duration of the signal is $2$ ms, the observer receives light from different regions of the screen with slightly varied arrival times, which results in a modest extension of the observed signal duration—an effect that becomes more pronounced at lower frequencies. This phenomenon has the same physical origin as scattering broadening, wherein variations in path length elongate the temporal profile of the signal.
@@
Furthermore, at lower frequencies the observer (red star) lies within the inner caustic region (as seen in Figures 15(a) and (b)), where little to no light is received. As the frequency increases, the caustics shift toward the center and, upon crossing the observer, sharp spikes appear. In the region between the two caustics, the intensity remains elevated relative to the non-lensed case, owing to the superposition of uniform and scattered light from the central region. Beyond the outer caustic, the signal reverts to its normal, unaltered state. In the spectral domain, these effects results in a characteristic two-spike feature, as predicted by Cordes et al. (2017). Additionally, interference fringes and multi-scale variations emerge. These complex features arise from the interference of numerous light rays with varying phases, with a weak beat-like effect—stemming from the interplay between the scintillation bandwidth and the spectral resolution—further modulating the pattern.
@@
The model also produces pronounced spikes on the standard scintillation pattern due to the redistribution of light by nonuniform plasma density variations. This phenomenon could potentially help understand the narrow-band properties. Large clumps, on the order of AU, can act as plasma lenses that alter the paths and intensities of light. While a single Gaussian lens typically generates a two-peak scintillation pattern in the spectrum, the presence of multiple clumps can lead to several spectral peaks at the observer. Unlike normal scintillation, where enhanced intensity results from constructive interference, these peaks arise from the convergence of redistributed light, thereby increasing the signal intensity at specific frequencies. An example of this effect is shown in Figure 17. The screen contains numerous large clumps, each approximately $1$ AU in size, with a 3D Gaussian density distribution and a central maximum density of $\sim 1\,\rm cm^{-3}$– characteristics similar to those depicted in Figure 3 (a). Each clump acts as a lens, producing a ring-like projection on the observer plane. At a given observer location, only a subset of these clumps effectively deflects light toward the observer, thereby contributing to the final waterfall plot. In this configuration, the screen is placed $50$ kpc from the observer plane and $1$ Gpc from the source, with the contributing clumps situated roughly 1 AU from the line connecting the source and observer, resulting in scintillation stripes. The prominent spikes in the received signal correspond to caustics produced by individual clumps.
@@ 5 Summary and Discussion
With the rapid advancements in both observations and theoretical studies in the FRB field, the demand for precise and comprehensive numerical methods to investigate propagation effects has significantly increased. This study presents a numerical framework, EMPI, to model the interaction between electromagnetic waves and plasma, specifically addressing linear, non-magnetized plasma-induced propagation effects on FRBs. The framework unifies all four types of non-magnetized plasma effects, enabling a comprehensive analysis of plasma-induced distortions in FRB signals. The functionality and reliability of the method are rigorously validated by comparing numerical results with analytical solutions or estimates. Additionally, we present examples involving more complex scenarios to demonstrate the capability of this framework in addressing realistic problems. The numerical results confirm the accuracy and robustness of the code, establishing EMPI as a reliable and valuable tool for studying plasma-induced propagation effects.
Although the initial results appear promising, EMPI remains a newly developed rudimentary code and has certain limitations: (i) As discussed in Section 2.1, selecting an appropriate computational regime is crucial. However, the current selection process lacks clarity and convenience, requiring researchers to align the regime with their specific objectives and evaluation criteria manually. To address this limitation, future updates will introduce more scientifically rigorous and quantitatively defined selection criteria. Additionally, a module will be developed to analyze plasma density grids, enabling automatic identification of plasma structures and regime selection based on newly established, more computationally applicable criteria. (ii) As noted in Section 4, some numerical results exhibit relatively large errors due to various factors. These issues emphasize the importance of careful parameter selection and appropriate resolution settings to optimize accuracy. Though detailed guidelines for standard outputs and code operation are provided, further optimizations could enhance user experience and computational efficiency. (iii) Reconstructing the intervening plasma distribution from observational data remains challenging. The complexity increases significantly in three-dimensional cases, where multiple parameters collectively contribute to the final received signals. Different parameter combinations can yield similar observational results, and for a given plasma distribution, different observer positions lead to varying observations. Nevertheless, this approach provides a means to constrain plasma parameters, particularly when incorporating the effects of magnetic fields and FRB temporal evolution. Additional observational information will help further refine these constraints, improving the accuracy of parameter estimation.
@@
Despite these limitations, EMPI offers substantial potential for future applications in plasma propagation studies. Several key directions include: (a) While most previous studies have focused on one-dimensional analytical plasma lenses, real astrophysical plasma distributions are inherently three-dimensional. The EMPI framework allows for detailed studies of complex plasma distributions. Even when the total column density remains constant, different spatial distributions can result in distinct propagation effects. In an extreme case where the column density follows a Gaussian distribution due to plasma spatial arrangement, some regions contain plasma while others do not. Within plasma-containing regions, the density remains uniform, but the plasma as a whole follows a Gaussian distribution in space. As a result, the overall column density appears Gaussian, yet no deflection occurs in uniform-density regions. The primary goal of this paper is to introduce a comprehensive numerical framework that facilitates the exploration of various plasma distributions, rather than to explain the underlying physics of specific distributions. More complex cases and discussions will be addressed in future work. (b) This approach bridges the gap between idealized analytical plasma models and fully turbulent distributions characterized by their statistical properties. By enabling more realistic numerical calculations, the EMPI framework provides a means to constrain FRB and plasma parameters by direct comparison with observations. The utilization of Markov Chain Monte Carlo (MCMC) methods with this framework will further refine constraints on the properties of the intervening plasma and FRBs. (c) The flexibility to adjust the positions of both the screen and the observer allows for the incorporation of velocity terms. This capability will enable calculations of scintillation arc phenomena using time-series snapshots, providing a new approach to studying FRB variability and propagation effects. (d) The methodology presented in this study is applicable to other scenarios where similar conditions hold. In particular, it can be naturally extended to other radio sources, especially transients. If the computational method criteria are met, this approach could also be extended to other wavelengths, broadening its applicability in astrophysics.
@@
In summary, this work contributes a robust methodology for analyzing electromagnetic wave propagation through plasma in astrophysical contexts, contributing to the growing body of research on plasma interactions. The effects of magnetic fields on polarization in magnetized plasma will be explored in the next paper of this series.
@@ 6 Acknowledgments
N.X., H.G., T.-C.W., and R.G. are supported by the National Natural Science Foundation of China (Projects 12373040,12021003), the National SKA Program of China (2022SKA0130100) and the Fundamental Research Funds for the Central Universities. Y.-P.Y. is supported by the National Natural Science Foundation of China grant No.12473047, the National Key Research and Development Program of China (2024YFA1611603) and the National SKA Program of China (2022SKA0130100). W.-Y. W. acknowledges support from the NSFC (No. 12261141690 and No. 12403058), the National SKA Program of China (No. 2020SKA0120100), and the Strategic Priority Research Program of the CAS (No. XDB0550300).

### v2

@@ 4.1.5 Convergence Test
The effect of resolution on the received signal is demonstrated in Figure 14. With a fixed plasma screen and observer position, the resolution is varied by adjusting the grid size from $1$ au down to $10^{-5}$ au. At coarser resolutions, the results exhibit significant instability and large fluctuations in the signal. Under such low-resolution conditions, some frequency components may exhibit detectable signal arrivals while others remain undetected, leading to significant stochastic variations. Further more, regions accessible to photon detection tend to exhibit uniform intensity characteristics due to the limited number of ray paths contributing to the composite signal.
As the resolution increases, the simulation results progressively stabilize. Finer grids yield waterfall plots with enhanced sharpness. Crucially, the spectral domain where caustic structures become localized near the observer displays pronounced intensity enhancement.
This stabilization confirms that resolution refinement enables precise representation of the light distribution and construction of the simulated waterfall plot, leading to a more faithful modeling of the authentic plasma lensing effect.
@@ 4.2 Extended Results
Expanding upon the foundational configurations established in preceding analyses, this section presents scenarios that incorporate conditions more representative of real-world astrophysical environments. These examples provide an opportunity to examine the applicability of this code to detailed astrophysical phenomena more closely, thereby showcasing its capability to handle intricate spatial configurations and variable parameters in complex real-world scenarios.
@@
Figure 15 displays the intensity distribution on the observer plane for a 2D Gaussian lens. The lens is modeled by the distribution
@@
which is analogous to the 1D Gaussian case (with $N_{\rm e}=n_{\rm e}l$, where $n_{\rm e}$, $l$, and $a$ defined as in the 1D formulation). In this example, it sets $N_{\rm 0}=0.015\,\rm pc\,cm^{-3}$, and $a=2$ au, with both the source-to-screen and screen-to-observer distances equal to $1$ kpc. The source is modeled as a point source with a half-open angle of $10^{-5}$ degrees, ensuring that the projected circle on the screen encompasses the entire lens. The color bar indicates the relative intensity normalized by the intensity in the absence of lensing.
Generally, the lens exhibits stronger effects at lower frequencies. Numerical calculations demonstrate bright annular regions (caustics) with dramatically enhanced relative intensities, precisely where theoretical treatments predict divergence. The lens behaves as a diverging element, redistributing light from the central region toward the periphery and thereby forming two distinct caustic rings. In regions distant from the lens, where the electron density is nearly uniform, no lensing occurs. This behavior is consistent with the findings of Clegg et al. (1998), which indicate that inner caustics are less sensitive to variations than outer caustics. The red star in Figure 15 marks the location of the observer, corresponding to the waterfall plots presented in Figure 16.
@@
Figure 16 illustrates the waterfall plots at the observer location marked by the filled red star in Figure 15. This plot highlights how the lensing structure alters the received signal, manifesting as variations in intensity and spectrum. The DM value of the plasma lens is sufficiently low and any background DM has been removed, resulting in a flat-top square wave pattern in the waterfall plot. Although the intrinsic duration of the signal is $2$ ms, the observer receives light from different regions of the screen with slightly varied arrival times, which results in a modest extension of the observed signal duration–an effect that becomes more pronounced at lower frequencies. This phenomenon has the same physical origin as scattering broadening, wherein variations in path length elongate the temporal profile of the signal.
@@
Furthermore, at lower frequencies, the observer (red star) lies within the inner caustic region (as seen in Figures 15(a) and (b)), where negligible light is received. As the frequency increases, the caustics shift toward the center and their intersection with the observer triggers abrupt intensity surges. In the annular region between the two caustics, the intensity remains elevated relative to the nonlensed case, owing to the superposition of uniform background intensity and scattered light from the central region. Exterior to the outer caustic, the signal reverts to its normal, unaltered state. In the spectral domain, these effects result in the characteristic two-spike feature, as predicted by Cordes et al. (2017). Additionally, interference fringes and multiscale variations emerge. These striped intricate patterns in the plot arise from the interference of numerous light rays with varying phases, with a weak beat-like effect–stemming from the interplay between the scintillation bandwidth and the spectral resolution–further modulating the pattern.
@@
The model generates distinct intensity spikes superimposed on the standard scintillation patterns through the redistribution of light caused by structured plasma density variations. This phenomenon could potentially help understand the narrowband properties of FRBs. Macroscopic clumps, on the order of an astronomical unit, can function as plasma lenses that alter the paths and received intensities of light. While a single Gaussian lens typically generates a two-peak scintillation pattern in the spectrum, the presence of multiple clumps can lead to several spectral peaks at the observer. Unlike normal scintillation, where enhanced intensity results from constructive interference, these peaks arise from the convergence of redistributed light, thereby increasing the signal intensity at specific frequencies. An example of this effect is shown in Figure 17. The screen contains numerous large clumps, each approximately $1$ au in size, with a 3D Gaussian density distribution and a central maximum density of $\sim 1\,\rm cm^{-3}$, characteristics similar to those depicted in Figure 3 (a). Each clump acts as a lens, producing a ring-like projection on the observer plane. At a given observer location, only a subset of these clumps effectively deflects light toward the observer, thereby contributing to the final waterfall plot. In this configuration, the screen is placed $50$ kpc from the observer plane and $1$ Gpc from the source, with the contributing clumps situated roughly $1$ au from the line connecting the source and observer, resulting in scintillation stripes. The prominent spikes in the received signal correspond to caustics produced by individual clumps.
@@ 5 Summary and Discussion
With the rapid advancements in both observations and theoretical studies in the FRB field, the demand for precise and comprehensive numerical methods to investigate propagation effects has significantly increased. This study presents a numerical framework, EMPI, to model the interaction between electromagnetic waves and plasma, specifically addressing linear, nonmagnetized plasma-induced propagation effects on FRBs. The framework unifies all four types of non-magnetized plasma effects, enabling a comprehensive analysis of plasma-induced distortions in FRB signals. The functionality and reliability of the method are rigorously validated by comparing numerical results with analytical solutions or estimates. Additionally, we present examples involving more complex scenarios to demonstrate the capability of this framework in addressing realistic problems. The numerical results confirm the accuracy and robustness of this code, establishing EMPI as a reliable and valuable tool for studying plasma-induced propagation effects.
As a nascent computational tool, EMPI currently exhibits three principal limitations requiring methodological refinement: (i) As discussed in Section 2.1, the regime selection mechanism lacks automated quantitative criteria, requiring researchers to align the regime with their specific objectives and evaluation criteria manually. To address this limitation, future updates will introduce more scientifically rigorous and quantitatively defined selection criteria. Additionally, a module will be developed to analyze plasma density grids, enabling automatic identification of plasma structures and regime selection based on newly established, more computationally applicable criteria. (ii) As noted in Section 4, parameter sensitivity exerts substantial influence on numerical outcomes. Benchmark studies demonstrate that the selection of critical parameters and grid resolution decisively governs solution accuracy. Although standardized output protocols and operational guide lines have been established, the implementation of advanced optimization algorithms remains imperative to enhance numerical stability. (iii) The reconstruction of intervening plasma distribution from observational data presents fundamental challenges. As propagation effects arise from the combined influence of multiple physical processes, inverting observational data to determine plasma configurations constitutes a significant scientific obstacle. In 3D cases, this complexity escalates substantially. Different parameter combinations can yield nearly identical observational signatures, and a fixed plasma distribution produces markedly distinct received signals when observed from varying positions. Nevertheless, this approach provides a viable framework for constraining plasma parameters. Future enhancements incorporating magnetic field effects and FRB temporal evolution characteristics will synergistically combine with additional multidimensional observational data, systematically improving parameter inversion accuracy through comprehensive constraint optimization.
@@
Notwithstanding these limitations, the EMPI framework demonstrates transformative potential in plasma propagation research through four key advancements: (a) Breakthrough in 3D structural effects. While prior studies predominantly employed idealized 1D analytic plasma lens models, EMPI pioneers high-resolution 3D plasma distribution calculation–critical for resolving parametric degeneracies in astrophysical plasmas. For instance, plasma configurations with identical column densities generate markedly different propagation signatures depending on their spatial morphology, a phenomenon quantitatively inaccessible through traditional analytic approaches. The primary goal of this paper is to introduce a comprehensive numerical framework that facilitates the exploration of various plasma distributions, rather than to explain the underlying physics of specific distributions. More complex cases and discussions will be addressed in future work. (b) Novel observational constraint methodology. This approach bridges the gap between idealized analytical plasma models and fully turbulent distributions characterized by their statistical proper ties. By enabling more realistic numerical calculations, the EMPI framework provides a means to constrain FRB and plasma parameters by direct comparison with observations. The utilization of Markov Chain Monte Carlo methods with this framework will further refine constraints on the properties of the intervening plasma and FRBs. (c) Dynamic evolution modeling capability. The flexibility to adjust the positions of both the screen and the observer allows for the incorporation of velocity terms. This capability will enable calculations of scintillation arc phenomena using time-series snapshots, providing a new approach to studying FRB variability and propagation effects. (d) Generalized astrophysical applicability. The methodology presented in this study is applicable to other scenarios where similar conditions hold. In particular, it can be naturally extended to other radio sources, especially transients. If the computational method criteria are met, this approach could also be extended to other wavelengths, broadening its applicability in astrophysics.
@@
In summary, this study establishes a robust methodology for analyzing electromagnetic wave propagation through astrophysical plasma, advancing the understanding of plasma mediated interactions. The subsequent paper in this series will focus on magnetized plasma environments, specifically analyzing how magnetic fields alter the polarization properties of electromagnetic signals.
@@ 6 Acknowledgments
N.X., H.G., T.-C.W., and R.G. are supported by the National SKA Program of China (2022SKA0130100), the National Natural Science Foundation of China (Projects 12373040,12021003) and the Fundamental Research Funds for the Central Universities. Y.-P.Y. is supported by the National Natural Science Foundation of China grant No.12473047, the National Key Research and Development Program of China (2024YFA1611603) and the National SKA Program of China (2022SKA0130100). W.-Y. W. acknowledges support from the NSFC (No. 12261141690 and No. 12403058), the National SKA Program of China (No. 2020SKA0120100), and the Strategic Priority Research Program of the CAS (No. XDB0550300).